1) It is friction that provides the force for a car to accelerate, so for high-performance cars the factor that limits acceleration isn't the engine; it's the tires.

For typical rubber-on-concrete friction, what is the shortest time in which a car could accelerate from 0 to 60 mph? Suppose that μs=1.00 and μk=0.80.

Express your answer to two significant figures and include the appropriate units.

2) A 22 cm diameter bowling ball has a terminal speed of 69 m/s. Suppose that the density of air is 1.2 kg/m3.

What is the ball's mass?

Express your answer to two significant figures and include the appropriate units.
Answer #1

Here ,

while accelerating as the tyres roll without slipping ,

maximum acceleration = 01 * 9.8

maximum acceleration = 9.8 m/s^2

initial speed , u = 0 m/s

final speed , v = 60 mph = 60 * 1609/3600

v = 26.8 m/s

Using first equation of motion

v = u + a * t

26.8 = 9.8 * t

t = 2.74 s

the time taken is 2.74 s